taylor co made a lump sum purchase of land, buildings, and equipment for 220,000. the appraised market value for each item is $20,000 land, $70,000, buildings, and 210,000 equipment. traulor should debit the equipment account for?
Answer #1

Total Appraised value = 20000+70000+210000 = 300000

Equipment = Lump sum value*Appraised value of equipment/Total Appraised value

= 220000*210000/300000

Equipment = 154000